OpenStudy (anonymous):

could someone help with this x/-7-9=-3

OpenStudy (anonymous):

x/-7 - 9 = -3 Add 9 to both sides. x/-7 = -3 + 9 x/-7 = 6 Multiply both sides by -7 x = -42